OpenStudy (anonymous):

For what values of x does the series below converge? What is its sum? What series do you get if you differentiate the given series term by term? For what values of x does the new series converge? What is its sum?

OpenStudy (anonymous):

\[1-\frac{ 1 }{ 2(x-3)}+\frac{ 1 }{ 4(x-3)^{2}}+...+(\frac{ -1 }{ 2 })^{n}(x-3)^{n}\]

OpenStudy (amistre64):

are you sure this is tped correctly

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

What I found... The series converges for 0<x<6

OpenStudy (anonymous):

The sum of the series within its interval of convergence is 3/x

OpenStudy (amistre64):

(x-3)^n when all the others have it in the denominator

OpenStudy (amistre64):

the series cant converge when x=3 since that would zero out the denominators

OpenStudy (anonymous):

oops. I wrote the question wrong.lol The (x-3) to the power of n are in the numerators.

OpenStudy (anonymous):

Sorry about that.

OpenStudy (amistre64):

a ratio test should help determine the interval of convergence

OpenStudy (anonymous):

Yeah I did do the ratio test and got 0<x<6

OpenStudy (amistre64):

lets dbl chk \[\lim_{n\to inf}\frac{(-2)^n(x-3)^{n+1}}{(-2)^{n+1}(x-3)^n}\] \[\lim_{n\to inf}\frac{x-3}{-2}\] \[\left|\frac{x-3}{-2}\right|\lim 1\] \[-1<\frac{x-3}{-2}<1\] \[-2<x-3<2\] \[1<x<5\]

OpenStudy (anonymous):

Oh okay, that makes sense

OpenStudy (anonymous):

but what about the abs signs? wouldn't it be +2 in denominator?

OpenStudy (anonymous):

oh wait nvm

OpenStudy (amistre64):

now, when x is between 1 and 5, the geometric ratio is less than 1 right? otherwise its more than 1

OpenStudy (anonymous):

yup! So then, the sum of the series would be 2/(x-1)?

OpenStudy (amistre64):

when x=1 or 5, the ratio is just an alternating addition of 1s which just bounces between 0 and 1

OpenStudy (amistre64):

well, r = (3-x)/2 for an infinite series with |r|<1 S = 1/(1-r) 1/(1-(3-x)/2) 2/(2-(3-x)) 2/(2-3+x) yes

OpenStudy (anonymous):

Okay, I found f'(x)= n(-1/2)^n (x-3)^(n-1) is that right?

OpenStudy (amistre64):

thats fair yes

OpenStudy (anonymous):

The series converges -2<x<3? not sure about this one..

OpenStudy (anonymous):

and im not sure how to find the sum for this one.

OpenStudy (amistre64):

brb ... 5 min tops

OpenStudy (anonymous):

lol okay :P

OpenStudy (amistre64):

had to do stuff ... its done now lol

OpenStudy (amistre64):

(n+1) (-1/2)^(n+1) (x-3)^((n+1)-1) ------------------------------- n (-1/2)^n (x-3)^(n-1) (n+1) (-1/2) ----------- n (x-3)^(-1) pull out all that has no n, and we are left with a lim of (n+1)/n which goes to 1 |3-x| < 2 is still our interval

OpenStudy (anonymous):

the same one as before, 1<x<5?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

so the sum would be the same?

OpenStudy (amistre64):

might want to work out a few terms, our constants are no longer constant are they

OpenStudy (anonymous):

Alright! Thanks!

OpenStudy (amistre64):

youre welcome

OpenStudy (amistre64):

S=1 +2r +3r^2 +4r^3 +5r^4+... -rS= -1r -2r^2 -3r^3 -4r^4-... ------------------------------ (1-r)S= 1 +r +r^2 +r^3 +r^4 + ... (1-r)S = (1-r^n)/(1-r) S = (1-r^n)/(1-r)^2 interesting

OpenStudy (amistre64):

does it work? lol

OpenStudy (anonymous):

lol idk.

OpenStudy (amistre64):

1/(1+(x-3)/2)^2 when x=2, the wolf says its 4 1/(1 -1/2)^2 1/(1/2)^2 = 4

OpenStudy (amistre64):

x=1.5 should be 16 by the wolf 1/ (1+(1.5-3)/2)^2 1/ ((2+1.5-3)/2)^2 4/ (2+1.5-3)^2 4/ (1/2)^2 = 16

OpenStudy (amistre64):

4/(x-1)^2 x=2 gives 4 x=1.5 ... thta should be it then

OpenStudy (amistre64):

lol, ive never tried that before ... that was fun :)

OpenStudy (anonymous):

lol ;p glad you enjoyed that :P